# Preparation of Papers for IEEE Sponsored Conferences & Symposia*

### v1

@@ IV Analyzing CCBS’s implemented branching rule
Since $j$ is stationary at $v(w_{j})$ throughout $t^{j},t^{j}+w^{j}_{D}]$ and the move action $m^{i}$ is identical in both $\langle m^{i},t^{i}\rangle$ and $\langle m^{i},t^{i}-\epsilon\rangle$, we can analyze when a collision occurs by comparing agent $i$’s motion in both cases. When agent $i$ executes $\langle m^{i},t^{i}-\epsilon\rangle$, its motion during the interval $t^{j},t^{j}+w^{j}_{D}-\epsilon]$ corresponds exactly to its motion during $t^{j}+\epsilon,t^{j}+w^{j}_{D}]$ when executing $\langle m^{i},t^{i}\rangle$. Since no collision occurred during $t^{j}+\epsilon,t^{j}+w^{j}_{D}]$ when executing $\langle m^{i},t^{i}\rangle$, and agent $j$ remains stationary at the same location $v(w^{j})$, no collision can occur during $t^{j},t^{j}+w^{j}_{D}-\epsilon]$ in the shifted case. Therefore, any collision between $i$ and $j$ executing $\langle m^{i},t^{i}-\epsilon\rangle$ and $\langle w^{j},t^{j}\rangle$ must occur during the remaining time interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$. ∎
Figure 5: Timeline of a conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}$ (red). The resulting constraints forbid $i$ from executing $m^{i}$ starting at $t^{i},t^{i}_{u})$ (orange) and $j$ from occupying $v(w^{j})$ during $I^{c}$ (purple). The timed action $\langle m^{i},t^{i}_{u}\rangle$ and two forbidden timed actions $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$ are shown.
Theorem IV.2 (Unsoundness of the implemented branching rule).
The implemented branching rule is not sound.
@@
Consider a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}=[t_{1}^{c},t_{2}^{c})$. From this, the implemented branching rule uses the constraint pair $\langle\langle i,m^{i},[t^{i},t^{i}_{u})\rangle,\langle j,v(w^{j}),I^{c}\rangle\rangle$. Letting $C_{i}\Leftrightarrow\langle i,m^{i},[t^{i},t^{i}_{u})\rangle$ and $C_{j}\Leftrightarrow\langle j,v(w^{j}),I^{c}\rangle$, the branching rule equivalently uses the constraint set pair $\langle C_{i},C_{j}\rangle$. Figure 5 illustrates a timeline of this conflict and the constrained intervals $t^{i},t^{i}_{u})$ and $I^{c}$.
By construction of $C_{i}$, we know that $c_{i}=\langle i,m^{i},t^{i}_{u}-\epsilon\rangle\in C_{i}$ for an arbitrarily small $\epsilon>0$. By definition of $t^{i}_{u}$, $\langle m^{i},t^{i}_{u}\rangle$ and $\langle w^{j},t^{j}\rangle$ do not conflict. It then follows by Lemma IV.1 that any collision between agent $i$ executing $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $j$ executing $\langle w^{j},t^{j}\rangle$ must occur during the time interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$. See Figure 5 for the timelines of the timed move actions $\langle m^{i},t^{i}_{u}\rangle$ and $\langle m^{i},t^{i}_{u}-\epsilon\rangle$.
The constraint set $C_{j}$ contains one forbidding constraint for each timed action that requires occupying $v(w^{j})$ during $I^{c}$. We now aim to find a forbidden timed wait action $\langle w^{\prime},t^{\prime}\rangle$ ($\langle j,w^{\prime},t^{\prime}\rangle\in C_{j}$) that does not conflict with $\langle m^{i},t^{i}_{u}-\epsilon\rangle$. Let
@@
that is, we restrict our search to timed wait actions that execute during a subset time interval of when $\langle w^{j},t^{j}\rangle$ is executed. Recall that a collision between $i$ and $j$ executing $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{j},t^{j}\rangle$ can only occur during $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$. Therefore, the same is true when executing $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$, a collision can only occur during the same interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$ (assuming $\epsilon<w^{\prime}_{D}$). From this, we now constrain $t^{\prime}+w^{\prime}_{D}<t^{j}+w^{j}_{D}-\epsilon$ such that $\langle w^{\prime},t^{\prime}\rangle$ and $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ do not collide:
@@
Finally, for $\langle j,w^{\prime},t^{\prime}\rangle\in C_{j}$ we require $t^{\prime},t^{\prime}+w^{\prime}_{D}]$ to intersect with $I^{c}$. Since $\epsilon$ is arbitrarily small, let $\epsilon<|I^{c}|$. It follows that the interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$ of size $\epsilon$ can only span at most a subset of $I^{c}$. Additionally, by (5), $I^{c}\subseteq[t^{j},t^{j}+w^{j}_{D}]$ such that $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$ must be a suffix of $I^{c}$. The interval $t_{1}^{c},t^{j}+w^{j}_{D}-\epsilon)$ is therefore non-degenerate. Thus, we require (6) and $t_{1}^{c}\leq t^{\prime}+w^{\prime}_{D}$ to hold. Since any $t^{\prime}\in\Re$ and $w^{j}_{D}\in\Re$ can be selected, we know that there exists a forbidding constraint $c_{j}=\langle j,w^{\prime},t^{\prime}\rangle\in C_{j}$. See Figure 5 for an example of such a timed wait action $\langle w^{\prime},t^{\prime}\rangle$.
@@
In conclusion, the timed actions $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$ do not conflict; $i$ collides with $j$ stationary at $v(w^{j})$ only at some time in $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$, while $j$ executing $\langle w^{\prime},t^{\prime}\rangle$ does not occupy $v(w^{j})$ at any time in $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$. Therefore, a solution $\Pi$ can contain both of these timed actions. However, such a solution does not satisfy $c_{i}(\Pi)\vee c_{j}(\Pi)$. Thus, $\langle c_{i},c_{j}\rangle$ is not sound, hence neither is $\langle C_{i},C_{j}\rangle$. By Lemma III.1, the implemented branching rule is not sound. ∎
V A Branching Rule for Sound and Solution-Complete CCBS
@@
In this section, we present our novel branching rule and prove that CCBS using this branching rule is sound and solution complete. To do so, we first show in Section V-A that our branching rule is sound. To provide termination guarantees, we formulate a non-termination condition in Section V-B and then show in Section V-C that CCBS under our branching rule does satisfy this condition. Finally, we use these results to conclude in Section V-D that CCBS under our branching rule is sound and solution complete.
@@
Our branching rule is based on shifting constraints [28] can be understood by considering a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$, with intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$. By construction of $\bar{I^{c}}$, when $i$ executes $\langle m^{i},t^{i}\rangle$ then a collision occurs if $j$ occupies $v(w^{j})$ at any time in $\bar{I^{c}}$. Suppose now that $i$ executes $\langle m^{i},t^{i}+\Delta t\rangle$ with $\Delta t\in\Re$ instead. Then, the intersection interval will also shift by $\Delta t$, such that a collision occurs if $j$ occupies $v(w^{j})$ at anytime in $\bar{t_{1}^{c}}+\Delta t,\bar{t_{2}^{c}}+\Delta t)$. Fig. 6 illustrates the timeline of $m^{i}$ starting at $t^{i}$ and incrementally later times. Now, let some $0\leq\delta<|\bar{I^{c}}|$ be given. If $i$ executes $\langle m^{i},t^{i}+\Delta t\rangle$ for any $0\leq\Delta t\leq\delta$, with shifted intersection interval $\bar{t_{1}^{c}}+\Delta t,\bar{t_{2}^{c}}+\Delta t)$, then we know for any value of $\Delta t$ that the interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ will be included in the resulting intersection interval. In other words, if $i$ executes $\langle m^{i},t^{i}+\Delta t\rangle$ for any $0\leq\Delta t\leq\delta$, then we know that a collision will arise if agent $j$ occupies $v(w^{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$. This result is useful as it allows us to form a pair of sound constraints from a move-wait conflict.
Figure 6: If move action $m^{i}$ is executed at any time in the (orange) interval $t^{i},t^{i}+\delta)$, then $i$ will necessarily intersect with $j$ waiting at $v(w)$ during the (blue) interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
We now define our branching rule.
Definition V.1 (Our novel branching rule).
Move-move and move-wait conflicts are handled differently:
@@
  • Given a move-move conflict $\langle\langle m^{i},t^{i}\rangle,\langle m^{j},t^{j}\rangle\rangle$, the theoretical branching rule is applied. That is, the unsafe intervals $t^{i},t^{i}_{u})$ and $t^{j},t^{j}_{u})$ are found. Agent $i$ is forbidden from executing $\langle m^{i},t\rangle$ for any $t\in[t^{i},t^{i}_{u})$, and $j$ is forbidden from executing $\langle m^{j},t\rangle$ for any $t\in[t^{j},t^{j}_{u})$.
@@
  • Given a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}=[t_{1}^{c},t_{2}^{c})$ and intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$, let
@@
    where $0<\gamma<1$ is a fixed constant. Agent $i$ is forbidden from executing $\langle m^{i},t\rangle$ for any $t\in[t^{i},t^{i}+\delta)$. Agent $j$ is forbidden from occupying the vertex $v(w_{j})$ during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ (vertex constraint) and executing any timed move action $\langle m,t\rangle$ where $\mathit{from}(m)=v(w^{j})$ and $t\in[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
The fixed constant $\gamma\in(0,1)$ dictates which agent in a move-wait conflict is more heavily constrained: setting $\gamma$ near $1$ increases the constrained interval on agent $i$ while setting it closer to 0 increases the constrained interval on $j$. Figure 7 illustrates three different cases where our branching rule is applied to a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$. In the first case $\delta=t_{2}^{c}-\bar{t_{1}^{c}}$ while in the second and third case $\delta=\gamma|\bar{I^{c}}|$.
Figure 7: $\delta$ for three different move-wait conflicts, with $\gamma=0.6$. In the top cases, $\delta=t_{2}^{c}-\bar{t_{1}^{c}}$, in the bottom two cases $\delta=\gamma(\bar{t_{2}^{c}}-\bar{t_{1}^{c}})$. Orange interval is the restriction on the move action, purple is the restriction on the wait action.
V-A Soundness
We now show that our branching rule is sound. The branching rule applies the same rule as in CCBS on move-move conflicts, which is shown in Section 4.2 of [20] to satisfy property (2) and is therefore sound. Hence, we refer the reader to there for more details. What remains is to prove the soundness of our branching rule on move-wait conflicts.
Theorem V.1 (Soundness of our branching rule).
Our branching rule is sound when applied to a move-wait conflict.

### v2

@@ 4 Analyzing CCBS’s Implemented Branching Rule
Since $j$ is stationary at $v(w_{j})$ throughout $t^{j},t^{j}+w^{j}_{D}]$ and the move action $m^{i}$ is identical in both $\langle m^{i},t^{i}\rangle$ and $\langle m^{i},t^{i}-\epsilon\rangle$, we can analyze when a collision occurs by comparing agent $i$’s motion in both cases. When agent $i$ executes $\langle m^{i},t^{i}-\epsilon\rangle$, its motion during the interval $t^{j},t^{j}+w^{j}_{D}-\epsilon]$ corresponds exactly to its motion during $t^{j}+\epsilon,t^{j}+w^{j}_{D}]$ when executing $\langle m^{i},t^{i}\rangle$. Since no collision occurred during $t^{j}+\epsilon,t^{j}+w^{j}_{D}]$ when executing $\langle m^{i},t^{i}\rangle$, and agent $j$ remains stationary at the same location $v(w^{j})$, no collision occurs during $t^{j},t^{j}+w^{j}_{D}-\epsilon]$ in the shifted case. Therefore, any collision between $i$ and $j$ executing $\langle m^{i},t^{i}-\epsilon\rangle$ and $\langle w^{j},t^{j}\rangle$ must occur during the remaining time interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$. ∎
Figure 5: Timeline of a conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with intersection interval, $\bar{I^{c}}$ and the collision interval in red. The resulting constraints forbid $i$ from executing $m^{i}$ starting at $t^{i},t^{i}_{u})$ (orange) and $j$ from occupying $v(w^{j})$ during $\bar{I^{c}}$ (purple). The timed action $\langle m^{i},t^{i}_{u}\rangle$ and two non-conflicting forbidden timed actions $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$ are shown.
Theorem 4.2 (Unsoundness of the IBR).
CCBS’s IBR is not sound.
@@
Consider a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$. From this, the IBR creates the constraints $\langle i,m^{i},[t^{i},t^{i}_{u})\rangle$ and $\langle j,v(w^{j}),\bar{I^{c}}\rangle$. Letting $C_{i}\Leftrightarrow\langle i,m^{i},[t^{i},t^{i}_{u})\rangle$ and $C_{j}\Leftrightarrow\langle j,v(w^{j}),\bar{I^{c}}\rangle$, the IBR branches using the constraint set pair $\langle C_{i},C_{j}\rangle$. The move-wait conflict and the constraint intervals $t^{i},t^{i}_{u})$ and $\bar{I^{c}}$ are illustrated in the upper part of Figure 5.
By construction of $C_{i}$, for $0<\epsilon<t^{i}_{u}-t^{i}$ we know that $\langle i,m^{i},t^{i}_{u}-\epsilon\rangle\in C_{i}$. By definition of $t^{i}_{u}$, $\langle m^{i},t^{i}_{u}\rangle$ and $\langle w^{j},t^{j}\rangle$ do not conflict. It then follows from Lemma 4.1 that any collision between agent $i$ executing $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $j$ executing $\langle w^{j},t^{j}\rangle$ must occur during the time interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$. The timelines of $\langle m^{i},t^{i}_{u}\rangle$ and $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ are illustrated in the lower part of Figure 5.
The constraint set $C_{j}$ contains one forbidding constraint for each timed action that requires occupying $v(w^{j})$ during $\bar{I^{c}}$. We now aim to find a forbidden timed wait action $\langle w^{\prime},t^{\prime}\rangle$ (i.e., $\langle j,w^{\prime},t^{\prime}\rangle\in C_{j}$) that does not conflict with $\langle m^{i},t^{i}_{u}-\epsilon\rangle$. Let
@@
that is, we restrict our search to timed wait actions that execute during a subset time interval of when $\langle w^{j},t^{j}\rangle$ is executed. Recall that a collision between $i$ and $j$ executing $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{j},t^{j}\rangle$ can only occur during $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$. The same also holds for $\langle w^{\prime},t^{\prime}\rangle$ since it entails $j$ waiting stationary at the same vertex; a collision between $i$ and $j$ executing $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$ can only occur during $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$. We can use this by constraining $\langle w^{\prime},t^{\prime}\rangle$ to end before the start of this interval, that is $t^{\prime}+w^{\prime}_{D}<t^{j}+w^{j}_{D}-\epsilon$, and by doing so ensuring that $\langle w^{\prime},t^{\prime}\rangle$ and $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ do not collide:
@@
Finally, to ensure that $\langle j,w^{\prime},t^{\prime}\rangle$ is indeed a member of $C_{j}$, we require $t^{\prime},t^{\prime}+w^{\prime}_{D}]$ to intersect with $\bar{I^{c}}$. Since $\epsilon$ can be arbitrarily small, let $\epsilon<|\bar{I^{c}}|$. It follows that the interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$ of size $\epsilon$ can only span at most a subset of $\bar{I^{c}}$. Thus, a non-degenerate subset of $t^{j},t^{j}+w^{j}_{D}-\epsilon)$ is also included in $\bar{I^{c}}$. Therefore, there exists a $\langle w^{\prime},t^{\prime}\rangle$ such that (6) is satisfied and $t^{\prime},t^{\prime}+w_{D}^{\prime})\cap\bar{I^{c}}\neq\varnothing$. These properties ensure that $\langle w^{\prime},t^{\prime}\rangle$ does not conflict with $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and that $\langle j,w^{\prime},t^{\prime}\rangle\in C_{j}$. An example of such a timed wait action is illustrated at the bottom of Figure 5.
@@
In conclusion, the timed actions $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$ do not conflict. Therefore, a solution $\Pi$ can contain both of these timed actions. However, letting $c_{i}=\langle i,m^{i},t^{i}_{u}-\epsilon\rangle$ and $c_{j}=\langle j,w^{\prime},t^{\prime}\rangle$, $c_{i}(\Pi)\vee c_{j}(\Pi)$ is not satisfied. Thus, $\langle c_{i},c_{j}\rangle$ is not sound, and since $c_{i}\in C_{i}$ and $c_{j}\in C_{j}$, $\langle C_{i},C_{j}\rangle$ is not sound either. By Lemma 3.1, the IBR is therefore not sound. ∎
5 A Branching Rule for Sound and Solution-Complete CCBS
@@
In this section, we present $\delta$-BR and prove that CCBS using $\delta$-BR is sound and solution complete. To do so, we first show in Section 5.1 that $\delta$-BR is sound. Termination guarantees are provided by formulating a non-termination condition in Section 5.2 and then showing in Section 5.3 that CCBS under $\delta$-BR does not satisfy this condition. Finally, these results are used in Section 5.4 to conclude that CCBS under $\delta$-BR is sound and solution complete.
@@
$\delta$-BR is based on shifting constraints [29] which can be understood by considering a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$. By definition of $\bar{I^{c}}$, when agent $i$ executes $\langle m^{i},t^{i}\rangle$ then a collision occurs if agent $j$ occupies $v(w^{j})$ at any time in $\bar{I^{c}}$. Suppose now that $i$ executes $\langle m^{i},t^{i}+\Delta t\rangle$ with $\Delta t\in\Re$ instead. Then, the intersection interval will also shift by $\Delta t$ such that a collision occurs if $j$ occupies $v(w^{j})$ at any time in $\bar{t_{1}^{c}}+\Delta t,\bar{t_{2}^{c}}+\Delta t)$. Figure 6 illustrates the timeline of $m^{i}$ starting at $t^{i}$ and incrementally later times. Now, let some $\delta$ be given such that $0\leq\delta<|\bar{I^{c}}|$. Then, if agent $i$ executes $\langle m^{i},t^{i}+\Delta t\rangle$ for any $0\leq\Delta t\leq\delta$, we know that the shifted intersection interval $\bar{t_{1}^{c}}+\Delta t,\bar{t_{2}^{c}}+\Delta t)$ will at least contain the interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ (shown in Figure 6). Therefore, we know that if agent $i$ executes $\langle m^{i},t^{i}+\Delta t\rangle$ for any $0\leq\Delta t\leq\delta$, then agent $j$ cannot occupy vertex $v(w^{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ without a collision occurring. This result is useful as it allows us to form a pair of sound constraints from a move-wait conflict.
Figure 6: Shifting the start of the timed move action $\langle m^{i},t^{i}\rangle$ also shifts the intersection interval $\bar{I^{c}}$. Executing $m^{i}$ at any time in the (orange) interval $t^{i},t^{i}+\delta)$ will result in a shifted intersection interval which necessarily includes the (purple) interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
Definition 5.1 ($\delta$-BR).
$\delta$-BR handles move-move and move-wait conflicts differently:
@@
  • Given a move-move conflict $\langle\langle m^{i},t^{i}\rangle,\langle m^{j},t^{j}\rangle\rangle$, the TBR/IBR is applied: the unsafe intervals $t^{i},t^{i}_{u})$ and $t^{j},t^{j}_{u})$ are found and the pair of motion constraints $\langle i,m^{i},[t^{i},t^{i}_{u})\rangle$ and $\langle j,m^{j},[t^{j},t^{j}_{u})\rangle$ are used to branch on. That is, agent $i$ is forbidden from executing $\langle m^{i},t\rangle$ for any $t\in[t^{i},t^{i}_{u})$ and agent $j$ is forbidden from executing $\langle m^{j},t\rangle$ for any $t\in[t^{j},t^{j}_{u})$.
@@
  • Given a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$, let
@@
    where $0<\gamma<1$ is a fixed constant. For agent $i$, the motion constraint $\langle i,m^{i},[t^{i},t^{i}+\delta)\rangle$ is created. For agent $j$, the vertex constraint $\langle j,v(w^{j}),[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ and motion constraints $\langle j,m,[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ for every $m\in\mathcal{A}:\mathit{from}(m)=v(w^{j})$ is created. That is, agent $i$ is forbidden from executing $m^{i}$ at any time in $t^{i},t^{i}+\delta)$, and agent $j$ is forbidden from occupying $v(w_{j})$ during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ or executing any move action from $v(w_{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
The fixed constant $\gamma\in(0,1)$ controls which agent in a move-wait conflict is more constrained: setting $\gamma$ near $1$ increases the constrained interval on the moving agent while setting $\gamma$ closer to 0 increases the constrained interval on waiting agent. Figure 7 illustrates four different cases where $\delta$-BR (with $\gamma=0.7$) is applied to a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$. In the first two cases (Figures 7(a) and 7(b)) we have $\delta=t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$ since $t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}<\gamma|\bar{I^{c}}|$. Thus, agent $j$ is forbidden from occupying $v(w^{j})$ from $t^{j}+w^{j}_{D}$ until the end of $\bar{I^{c}}$. On the other hand, in the second two cases (Figures 7(c) and 7(d)) we have $\delta=\gamma|\bar{I^{c}}|$. Thus, a non-singular part of agent $j$’s occupation of $v(w^{j})$ is included in the constrained interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$. Furthermore, it is also visible that agent $j$’s constrained interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ is smaller than agent $i$’s constrained interval $t^{i},t^{i}+\delta)$ as a result of $\gamma>0.5$.
(a) $\delta=t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$
(b) $\delta=t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$
(c) $\delta=\gamma|\bar{I^{c}}|$
(d) $\delta=\gamma|\bar{I^{c}}|$
Figure 7: An illustration of $\delta$-BR applied to four different move-wait conflicts, $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$, where the start time $t^{j}$ of wait action $w^{j}$ progressively increases for each sub-figure. $\delta=\min(\gamma|\bar{I^{c}}|,\;t^{j}+w^{j}_{D}-\bar{t_{1}^{c}})$ with $\gamma=0.7$ is shown; $\delta=t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$ in Figures 7(a) and 7(b), and $\delta=\gamma|\bar{I^{c}}|$ in Figures 7(c) and 7(d). Agent $i$’s constrained interval $t^{i},t^{i}+\delta)$ (orange) and agent $j$’s constrained interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ (purple) are shown.
5.1 Soundness
We now show that $\delta$-BR is sound. $\delta$-BR applies TBR/IBR on move-move conflicts, which is shown in Section 4.2 of [21] to satisfy property (2) and is therefore sound. Hence, we refer the reader to there for more details. What remains is to prove the soundness of $\delta$-BR on move-wait conflicts.
Theorem 5.1 (Soundness of $\delta$-BR).
$\delta$-BR is sound when applied to a move-wait conflict.